# An Efficient Instance Segmentation Framework Based on Oriented Bounding Boxes

### v3

@@ IV-E3 Oriented Bounding Box Detection
To explore the impact of the OBB detection module on the instance segmentation performance, we compare Oriented R-CNN used in CFNet with some other OBB detection methods [26, 27]. In addition, a HBB detection method (HBB detection branch of Mask R-CNN) is also added for comparison. As shown in Table VI, the effective use of OBBs significantly improves instance segmentation accuracy, and CFNet is not sensitive to specific OBB detection methods.
TABLE VI: Performance comparison of OBB and HBB detectors.
Mask Oriented Oriented S2ANet [27]
R-CNN [16] R-CNN [14] Reppoints [26]
42.95 64.29 64.37 64.06
V Conclusions
This paper proposes CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation in robot vision measurement. CFNet uses BSMs and introduces a novel OBB prompt encoder. To make CFNet more lightweight, a knowledge distillation method involving Gaussian smoothing for teacher model outputs is applied to the OBB prompt encoder and mask decoder. Experiments on both industrial and public datasets demonstrate that CFNet outperforms current instance segmentation methods.
One limitation of this work is that CFNet relies on prior geometric properties to achieve completely occluded object instance segmentation. In the future, we plan to utilize more types of prior knowledge to improve the robustness. It is worth noting that CFNet not only uniformly handles the two difficulties of completely occluded and dense object instance segmentation, but also can be used for general instance segmentation tasks. Moreover, CFNet can also handle objects that are simultaneously in completely occluded and dense environments, such as instance segmentation of containers with UAVs. We also plan to explore more applications of CFNet in various complex robotic tasks.

### v6

@@ IV-D3 Oriented Bounding Box Detection
To explore the impact of the OBB detection module on the instance segmentation performance, we compare Oriented R-CNN used in OBSeg with some other OBB detection methods, such as Oriented Reppoints [53] and S2ANet [52]. In addition, the HBB detection method YOLOv7 is also added for fair comparison. As shown in Table VII, the effective use of OBBs significantly improves instance segmentation accuracy, and OBSeg is not sensitive to specific OBB detection methods.
V Discussion
V-1 Comparison with Other Methods Using BSMs
Currently, there are some works that use BSMs, such as SAM, for instance segmentation tasks. These methods mainly aim to leverage the powerful segmentation capabilities of BSMs, which are acquired after training on large-scale datasets. However, these methods rely on HBB prompt input, which introduces many interference regions when performing instance segmentation on objects densely packed in multiple orientations, thereby reducing the segmentation performance. Different from these methods, our proposed OBSeg uses more accurate OBBs as prompts and achieves better instance segmentation performance.
Moreover, we introduce Gaussian smoothing-based knowledge distillation to make OBSeg more lightweight. Experimental results on iSAID, NWPU VHR-10, and PSeg-SSDD datasets demonstrate that OBSeg is superior to current BSM-based methods in terms of instance segmentation accuracy and has competitive inference speed.
V-2 Analysis of Inference Accuracy and Speed
Compared with methods using HBBs or those not based on BSMs, OBSeg uses BSMs with OBB prompts, which have fewer interference regions and are less dependent on bounding box detection performance. This enables OBSeg to achieve higher instance segmentation accuracy on multiple datasets.
In addition, since BSMs have powerful segmentation capabilities but slow inference speed, we have applied model lightweighting. To maintain high segmentation performance, we propose a knowledge distillation method based on Gaussian smoothing. Compared with existing methods, the lightweight distilled OBSeg has competitive inference speed while maintaining high segmentation accuracy.
V-3 Limitations and Future Work
Although OBSeg achieves better instance segmentation performance than existing methods, it also has some limitations. One limitation of this work is that the instance segmentation performance of OBSeg is not robust to objects with very large aspect ratios or extreme scales. This is mainly because these types of objects are in the minority in the dataset, and the model has insufficient feature recognition capabilities for objects with these characteristics. To address this type of long-tail distribution problem, future work will build upon the BSM framework by adopting class-balanced sampling strategies, extreme-scale feature learning, and adaptive adjustment of loss functions.
VI Conclusions
This paper proposes OBSeg, an accurate and fast instance segmentation framework using BSMs with OBB prompts in remote sensing images. OBSeg introduces a novel OBB prompt encoder, and we perform knowledge distillation to make OBSeg more lightweight and practical. Specifically, a Gaussian smoothing-based distillation method is introduced to improve the instance segmentation performance of lightweight distilled BSMs. Thanks to OBB prompts, OBSeg is less dependent on bounding box detection performance and has fewer interference regions than instance segmentation methods using HBBs or HBB prompts. Experiments on multiple public datasets, such as iSAID and NWPU VHR-10 datasets, demonstrate that OBSeg outperforms current instance segmentation methods in terms of instance segmentation accuracy and has competitive inference speed. One limitation of this work is that the instance segmentation performance of OBSeg is not robust to objects with very large aspect ratios or extreme scales. Future work will focus on improvements in data, extreme-scale feature learning, and loss functions.